# Adaptation of Agentic AI

### v2

@@ 5.2.2 Subagent-as-Tool
The year 2025 marked a paradigm shift in T2 research: the transition from training reactive tools (retrievers that respond to queries) to training proactive sub-agents (autonomous systems that actively explore, plan, orchestrate, and refine their operations over multiple turns while serving frozen primary agents). This shift was enabled by breakthroughs in reinforcement learning with verifiable rewards (RLVR) and represents the full maturation of the T2 vision: specialized sub-agents trained using frozen agent outputs as supervision signals, creating a symbiotic ecosystem where lightweight agentic tools co-evolve with frozen reasoning cores. Critically, this agentic transformation applies not only to information retrieval—training searchers that iteratively gather evidence—but also to meta-cognitive processes such as workflow orchestration and memory management, where sub-agents learn to coordinate frozen specialist modules and manage multi-step reasoning trajectories. We organize this evolution into three families of subagents: (i) agentic searchers, (ii) memory-construction subagents, (iii) meta-cognitive planners and orchestrators, and (iv) self-evolving subagents.
Agentic Searcher Breakthrough
s3 [27] (EMNLP 2025) demonstrated that training agentic tools could be radically more data-efficient than training agentic LLMs. The system trains a lightweight 7B “searcher” that performs multi-turn iterative search: generate queries, retrieve documents, select evidence, decide whether to search again or feed context to the frozen generator. The frozen generator (agent, Qwen2.5-14B or Claude) never updates, but provides the ultimate training signal through a metric called Gain Beyond RAG (GBR):
@@
rewarding the searcher only when the frozen model can answer correctly using its retrieved documents. This decoupling forces the subagent to optimize for retrieval quality rather than its own myopic behavior, reinforcing a core T2 principle: the frozen generator should not only consume a tool’s outputs but also supervise its learning.
Learning to Construct Memory as a Subagent
Extending beyond search, a complementary line of work treats long-term memory construction itself as a T2-style subagent problem. Mem-$\alpha$ [162] formulates memory management as RL over an explicit memory API, training a lightweight Qwen3-4B controller to operate a three-part external memory (core summary, semantic facts, episodic events) for a frozen backend generator. Only the memory-writing policy is optimized; the generator and retriever for downstream QA remain frozen. Rewards derive from verifiable outcomes—question-answering accuracy over long horizons, tool-call correctness, effective compression, and semantic validity of memory entries—so the subagent learns to construct compact yet sufficient memories that maximize the frozen model’s utility. Empirically, Mem-$\alpha$ significantly outperforms prior memory baselines and generalizes from $\sim$30k-token training sequences to contexts exceeding 400k tokens, exemplifying a memory-construction subagent that adaptively curates the information diet for a fixed reasoning core.
AutoGraph-R1 [163] applies this symbiotic principle to the construction of structured Knowledge Graphs (KGs). Rather than relying on static extraction heuristics, it optimizes an LLM-based “constructor subagent” to generate KGs from raw text. The supervision signal is derived directly from the frozen agent’s performance on downstream reasoning tasks (GraphRAG [164]) using the generated graph. This allows the constructor to learn a policy that prioritizes functional utility, creating connectivity and paths that specifically facilitate the host agent’s retrieval and reasoning, over intrinsic metrics like triple density.
Meta-cognitive and Control Subagents
Stepping further up the stack, recent work trains subagents that shape how frozen models think (planning, steering, and budgeting computation) rather than what they retrieve or store.
@@
where $R_{\text{outcome}}$ measures final task success, $R_{\text{process}}$ evaluates the rationality of the search plan (critiqued by the frozen generator), and Cost penalizes excessive planning. By leveraging both outcome and process rewards [165], the frozen model acts as executor and teacher, enabling the planner to internalize not only “what works” but “why it works.” Varying $\lambda$ traces a Pareto frontier between cost and quality, yielding planners tailored to different deployment budgets.
@@
Bridging from advising to driving, Matryoshka Pilot [167] (NeurIPS 2025) formalizes a controller–generator loop where a small white-box LLM controls a larger black-box LLM by emitting intermediate decomposition steps, plans, and summaries. Treating the black-box model as an environment, M-Pilot collects trajectory-level success signals and optimizes the controller with Iterative DPO. This yields $\approx$3–7% gains across reasoning, planning, and personalization benchmarks, and transfers plug-and-play across multiple black-box backends, further reinforcing the view of control subagents as portable T2 tools.
Learning to Orchestrate Frozen Specialists
@@
Orchestration-focused subagents push the T2 vision to its logical conclusion: training a dedicated policy to coordinate multiple frozen specialists.
@@
AgentFlow [51] decomposes an agent into modules—a planner, tool executor, verifier, and solution generator—implemented mostly as frozen Qwen2.5-7B-Instruct models. Only the planner is trained. Using Flow-GRPO, a single trajectory-level reward (correct vs. incorrect, judged by GPT-4o) is broadcast to all decisions in each rollout, with group-normalized advantages enabling effective credit assignment despite sparse rewards. Empirically, a 7B AgentFlow planner achieves 57.3% on search-intensive tasks (+14.9% over AutoGen), 51.5% on mathematical reasoning (+14.5% over ToRL), and 33.1% on GAIA—outperforming GPT-4 ($\sim$200B parameters) on several setups. This shows that learned orchestration of frozen specialists can rival or surpass monolithic models.
Self-Evolving (Sub)Agent
@@
A more advanced branch of the subagent-as-tool paradigm allows the tools themselves to co-evolve through self-generated tasks and rewards. R-Zero [168] instantiates two roles—a Solver and a Challenger—from the same base LLM. When the Solver is frozen, its successes, failures, and uncertainty (via self-consistency) define rewards that train the Challenger to propose tasks near the Solver’s capability frontier. Alternating these phases creates a bidirectional loop, yet each step still follows the T2 principle of optimizing a lightweight subagent under signals from a stronger or temporarily fixed core. Multi-Agent Evolve (MAE) [169] extends this design into a triadic architecture with a Proposer, Solver, and Judge. The Proposer and Judge operate as adaptive T2 subagents: the Judge learns to evaluate trajectories produced by the system, and the Proposer learns to generate diverse, high-quality, Solver-challenging tasks. Rather than tuning the main Solver, MAE improves performance by training these peripheral subagents to shape data, rewards, and curricula. Together, R-Zero and MAE illustrate a second generation of subagent-as-tool methods—self-evolving ecosystems that autonomously construct the learning conditions for otherwise frozen reasoning cores.
Synthesis: The Maturation of T2
@@
Across these lines of work, the subagent-as-tool paradigm extends T2 from what to retrieve, to what to remember, to how to plan, steer, and orchestrate, and finally to how to self-evolve. Agentic searchers (s3, DynamicRAG, QAgent) optimize information acquisition for frozen generators; memory-construction subagents (Mem-$\alpha$) curate long-horizon state; meta-cognitive controllers and orchestrators (AI-SearchPlanner, Advisor Models, Matryoshka Pilot, AgentFlow) decide how tools and specialists are deployed; and self-evolving frameworks (R-Zero, Multi-Agent Evolve) autonomously generate curricula and reward signals that continually refine these ecosystems. The consistent lesson is that decoupling tool training from generator training, while enabling tools to adapt to one another, yields systems that are more data-efficient, modular, generalizable, and robust than monolithic alternatives. In this mature T2 view, intelligence emerges not from scaling a single model, but from the learned coordination and co-evolution of specialized, frozen components through lightweight, adaptive subagents.
5.2.3 Agentic Memory and Others
An agent’s memory system itself can also be framed as an adaptive tool. Instead of modifying the agent’s core parameters, T2 methods can train or “tune” the memory module—how it writes, retrieves, reflects, and forgets—using the frozen agent’s downstream task performance or outputs as the supervisory signal. Survey on agent memory [48] categorizes a wide array of mechanisms that can be optimized as T2 tools. These works explore memory as a foundational component for agentic behavior, spanning short-term buffers, long-term experiential databases, and structured knowledge.
Dynamic Memory Stores
Many foundational memory architectures function as T2 tools, where the frozen agent’s output (e.g., a "write" command or new information) dynamically "tunes" the external memory store. This includes systems that manage short-term (in-context) and long-term (retrieval-based) storage, demonstrating how agents can manage context, retrieve past interactions, and maintain coherence over time by updating a peripheral memory tool [170, 171, 172, 173, 174, 175].
Experiential and Reflective Memory
A significant line of T2-aligned research focuses on memory modules that learn from experience. These tools enable a frozen agent to store, reflect on, and learn from its own output (e.g., entire trajectories), often using verbal reinforcement or self-correction. This allows the agent to build a curriculum of skills and avoid repeating past failures by tuning the memory tool without updating the core LLM’s weights [34, 176, 177, 178, 179, 180].
Structured Memory (Graphs, Trees, and Databases)
To move beyond linear text, some T2 memory tools structure information in sophisticated forms, such as knowledge graphs, trees, or symbolic databases. The frozen agent’s outputs are used as signals to “tune” this structured tool—for example, by adding new nodes, updating relationships, or writing to a database. These structured representations can be more efficiently queried and reasoned over by the frozen agent, effectively externalizing complex memory management into a specialized, adaptive tool [181, 182, 183, 184].
Episodic Memory as a Trainable Module

### v3

@@ 5.2.2 Subagent-as-Tool
The year 2025 marked a transition in T2 research from training reactive tools (retrievers that respond to queries) to training proactive subagents (autonomous systems that explore, plan, orchestrate, and refine their operations over multiple turns while serving frozen primary agents). The shift, enabled by advances in RLVR, applies not only to information retrieval but also to meta-cognitive processes such as workflow orchestration and memory management. The evolution can be organized into four families of subagents: (i) agentic searchers, (ii) memory-construction subagents, (iii) meta-cognitive planners and orchestrators, and (iv) self-evolving subagents.
Agentic searchers.
s3 [28] (EMNLP 2025) showed that training agentic tools can be radically more data-efficient than training agentic LLMs. The system trains a lightweight 7B “searcher” that performs multi-turn iterative search: generate queries, retrieve documents, select evidence, decide whether to search again or feed context to the frozen generator. The frozen generator (agent, Qwen2.5-14B or Claude) never updates, but provides the ultimate training signal through a metric called Gain Beyond RAG (GBR):
@@
rewarding the searcher only when the frozen model can answer correctly using its retrieved documents. The decoupling forces the subagent to optimize for retrieval quality rather than its own myopic behavior, reinforcing a core T2 principle: the frozen generator should not only consume a tool’s outputs but also supervise its learning.
Learning to construct memory as a subagent.
Extending beyond search, a complementary line of work treats long-term memory construction itself as a T2-style subagent problem. Mem-$\alpha$ [163] formulates memory management as RL over an explicit memory API, training a lightweight Qwen3-4B controller to operate a three-part external memory (core summary, semantic facts, episodic events) for a frozen backend generator. Only the memory-writing policy is optimized; the generator and retriever for downstream QA remain frozen. Rewards derive from verifiable outcomes (question-answering accuracy over long horizons, tool-call correctness, effective compression, and semantic validity of memory entries), so the subagent learns to construct compact yet sufficient memories that maximize the frozen model’s utility. In experiments, Mem-$\alpha$ outperforms prior memory baselines and generalizes from $\sim$30k-token training sequences to contexts exceeding 400k tokens, serving as a concrete instance of a memory-construction subagent that adaptively curates the information diet for a fixed reasoning core.
AutoGraph-R1 [164] applies this symbiotic principle to the construction of structured Knowledge Graphs (KGs). Rather than relying on static extraction heuristics, it optimizes an LLM-based “constructor subagent” to generate KGs from raw text. The supervision signal is derived directly from the frozen agent’s performance on downstream reasoning tasks (GraphRAG [165]) using the generated graph. The constructor thereby learns a policy that prioritizes functional utility, creating connectivity and paths that specifically facilitate the host agent’s retrieval and reasoning, over intrinsic metrics like triple density.
Meta-cognitive and control subagents.
Recent work trains subagents that shape how frozen models think (planning, steering, and budgeting computation) rather than what they retrieve or store.
@@
where $R_{\text{outcome}}$ measures final task success, $R_{\text{process}}$ evaluates the rationality of the search plan (critiqued by the frozen generator), and Cost penalizes excessive planning. By combining both outcome and process rewards [166], the frozen model acts as executor and teacher, so the planner internalizes not only “what works” but “why it works.” Varying $\lambda$ traces a Pareto frontier between cost and quality, yielding planners tailored to different deployment budgets.
@@
Bridging from advising to driving, Matryoshka Pilot [168] (NeurIPS 2025) formalizes a controller-generator loop where a small white-box LLM controls a larger black-box LLM by emitting intermediate decomposition steps, plans, and summaries. Treating the black-box model as an environment, M-Pilot collects trajectory-level success signals and optimizes the controller with Iterative DPO. The method yields $\approx$3–7% gains across reasoning, planning, and personalization benchmarks, and transfers plug-and-play across multiple black-box backends, further reinforcing the view of control subagents as portable T2 tools.
Learning to orchestrate frozen specialists.
@@
Orchestration-focused subagents train a dedicated policy to coordinate multiple frozen specialists.
@@
AgentFlow [52] decomposes an agent into modules (planner, tool executor, verifier, and solution generator) implemented mostly as frozen Qwen2.5-7B-Instruct models. Only the planner is trained. Using Flow-GRPO, a single trajectory-level reward (correct vs. incorrect, judged by GPT-4o) is broadcast to all decisions in each rollout, with group-normalized advantages permitting effective credit assignment despite sparse rewards. A 7B AgentFlow planner achieves 57.3% on search-intensive tasks (+14.9% over AutoGen), 51.5% on mathematical reasoning (+14.5% over ToRL), and 33.1% on GAIA, outperforming the much larger GPT-4 on several setups, demonstrating that learned orchestration of frozen specialists can rival or surpass monolithic models.
Self-evolving (sub)agent.
@@
A more advanced branch of the subagent-as-tool paradigm allows the tools themselves to co-evolve through self-generated tasks and rewards. R-Zero [169] instantiates two roles, a Solver and a Challenger, from the same base LLM. When the Solver is frozen, its successes, failures, and uncertainty (via self-consistency) define rewards that train the Challenger to propose tasks near the Solver’s capability frontier. Alternating these phases creates a bidirectional loop, yet each step still follows the T2 principle of optimizing a lightweight subagent under signals from a stronger or temporarily fixed core. Multi-Agent Evolve (MAE) [170] extends this design into a triadic architecture with a Proposer, Solver, and Judge. The Proposer and Judge operate as adaptive T2 subagents: the Judge learns to evaluate trajectories produced by the system, and the Proposer learns to generate diverse, high-quality, Solver-challenging tasks. Rather than tuning the main Solver, MAE improves performance by training these peripheral subagents to shape data, rewards, and curricula. Together, R-Zero and MAE illustrate a second generation of subagent-as-tool methods: self-evolving ecosystems that autonomously construct the learning conditions for otherwise frozen reasoning cores.
Synthesis: the maturation of T2.
@@
The subagent-as-tool paradigm progressively broadens the scope of T2: retrieval, memory construction, planning and orchestration, and finally self-evolution. Agentic searchers (s3, DynamicRAG, QAgent) optimize information acquisition for frozen generators; memory-construction subagents (Mem-$\alpha$) curate long-horizon state; meta-cognitive controllers and orchestrators (AI-SearchPlanner, Advisor Models, Matryoshka Pilot, AgentFlow) decide how tools and specialists are deployed; and self-evolving frameworks (R-Zero, Multi-Agent Evolve) autonomously generate curricula and reward signals. Decoupling tool training from generator training, while allowing tools to adapt to one another, yields systems that are more data-efficient, modular, and generalizable than monolithic alternatives.
5.2.3 Agentic Memory and Skills
An agent’s memory system can be framed as an adaptive tool. As discussed in §3, the paradigm label depends on the memory’s form and update mechanism: external non-parametric stores updated by a frozen agent’s outputs are predominantly T2; pre-trained or plug-in modules are T1; and parametric or hybrid architectures occupy the boundary between tool and agent adaptation. The majority of systems reviewed below are T2, where the frozen agent’s downstream task performance or outputs serve as the supervisory signal for tuning the memory module (how it writes, retrieves, reflects, and forgets). Several recent surveys provide complementary perspectives on this space. Hu et al. [171] propose a unified forms, functions, dynamics framework that organizes agent memory along three orthogonal dimensions: forms (token-level, parametric, and latent memory, distinguished by where and how information is stored), functions (factual memory for knowledge of users and environments, experiential memory for accumulated interaction outcomes, and working memory for active context management), and dynamics (the lifecycle of memory formation, evolution, and retrieval). This tripartite lens clarifies that memory adaptation operates simultaneously on the storage substrate, the content type, and the update policy. Within the experiential function, Hu et al. further distinguish three abstraction levels: case-based memory (raw trajectory storage), strategy-based memory (distilled workflows and heuristics), and skill-based memory (executable code, APIs, and MCP protocols). This hierarchy maps naturally onto our adaptation framework: case-based and strategy-based memory are predominantly T2 (the frozen agent’s success or failure determines what is retained), while skill-based memory bridges T2 (skill libraries curated by agent feedback) and T1 (pre-trained tool APIs that are agent-agnostic).
Zhang et al. [49] categorize the mechanisms that can be optimized as T2 tools, spanning short-term buffers, long-term experiential databases, and structured knowledge. Zhang et al. [172] provide a complementary survey focused specifically on the memory mechanisms of LLM-based agents, organizing existing work along the dimensions of memory formation (how memories are created), management (how they are updated and pruned), and utilization (how they inform future actions). Huang et al. [173] offer a broader treatment that connects cognitive-science memory models (complementary learning systems, working memory) to foundation-agent architectures, while Jiang et al. [174] provide an empirical analysis of evaluation and system limitations, identifying benchmark saturation, metric validity gaps, and backbone-model dependence as open problems. Jiang et al. [175] argue that long-term memory is the foundation of AI self-evolution, framing persistent memory as the substrate that enables agents to accumulate knowledge, refine strategies, and improve autonomously over extended horizons. At a higher level of abstraction, Sumers et al. [176] propose the Cognitive Architectures for Language Agents (CoALA) framework, which organizes agent memory into working memory (the active context window), episodic memory (records of past interactions), semantic memory (general knowledge), and procedural memory (learned action routines). CoALA provides a principled decomposition that maps directly onto the T2 design space: each memory type can be independently adapted under frozen-agent supervision, and the framework clarifies which memory subsystem should be optimized for a given task profile. Together, these surveys establish that agentic memory is not a single mechanism but a design space with multiple axes: storage modality (parametric vs. explicit vs. latent), temporal scope (transient, session-level, persistent), content type (factual, experiential, working), abstraction level (cases, strategies, skills), and update policy (append-only, reflective, RL-optimized).
Closely related is the concept of agent skills. Wu and Zhang [14] argue that agent skills are best understood through the lens of procedural memory: structured, reusable knowledge of how to perform tasks, paralleling the cognitive-science distinction between declarative knowledge (knowing that) and procedural knowledge (knowing how). Under this framing, skills follow a lifecycle of acquisition (learning from demonstrations, exploration, or trajectory distillation), representation (as executable code, API specifications, MCP protocols, or textual procedures), invocation (retrieval and execution at inference time), and refinement (iterative improvement through reflection, RL, or collective sharing). Xu and Yan [177] complement this procedural-memory view with a systems perspective: they emphasize progressive context loading, the role of SKILL.md and MCP as interface standards, and the security/governance problems that arise once skills become executable, shareable artifacts rather than passive memories. Fang et al. [178] provide an empirical investigation of procedural memory in LLM agents, showing that the format and granularity of stored procedures significantly affect downstream task performance, and that agents benefit from adaptive selection between code-based and natural-language procedure representations depending on task complexity. Cao et al. [179] propose a dynamic procedural memory framework in which the agent continuously refines its stored procedures based on execution outcomes, demonstrating that iterative memory refinement outperforms static memory accumulation on long-horizon planning tasks. The skill lifecycle maps onto the experiential memory hierarchy proposed by Hu et al. [171]: case-based memory stores raw trajectories, strategy-based memory distills transferable heuristics and workflows, and skill-based memory compiles executable capabilities. Memory and skills are thus two facets of the same adaptation mechanism. Memory provides the storage substrate and organizational structure; skills provide the executable, composable content that makes that storage actionable for future tasks.
Dynamic memory stores.
Foundational T2 memory architectures cluster into three design families based on how they organize stored information[180, 181, 182, 183, 184, 185]: hierarchical/OS-inspired systems (MemGPT, Memory OS) impose explicit tiers (working memory vs. long-term storage) with page-eviction or garbage-collection policies; reflection-based systems (Generative Agents, A-MEM) let the agent’s own outputs decide what to consolidate and when; and graph/structured systems (HippoRAG, SHIMI) index memories by relational or semantic structure rather than recency. The design choice determines the system’s trade-off between recall speed and associative richness. MemGPT [181] formalizes this idea through an operating-system inspired memory hierarchy: a limited main-context window acts as “RAM,” while an unbounded external store serves as “disk.” The agent issues explicit read/write operations to move information between tiers, and a page-eviction policy decides what to retain in the active context. Generative Agents [180] maintain a memory stream of natural-language observations and use the agent’s own reflections to periodically consolidate raw entries into higher-level abstractions, creating a two-tier store (observations and reflections) that supports planning over day-long time horizons. More recently, Gutiérrez et al. [186] recast the transition from static retrieval-augmented generation to persistent agent memory as a form of non-parametric continual learning, showing that memory stores can absorb new information without catastrophic forgetting of earlier knowledge, a property that parametric fine-tuning struggles to guarantee. Kang et al. [187] propose Memory OS, a layered architecture that treats memory management as an operating-system service, providing standardized APIs for storage, indexing, retrieval, and garbage collection that decouple memory logic from the agent’s reasoning loop. A-MEM [188] takes an agentic approach to memory management: the LLM itself decides how to organize its memory store, dynamically creating, linking, and restructuring entries based on content relevance and task demands rather than relying on fixed indexing heuristics. HippoRAG [189] draws on the neuroscience of hippocampal memory indexing, separating a neocortical component (an LLM that encodes passages into a knowledge graph) from a hippocampal index (a retrieval module that performs pattern completion over the graph). The design mirrors complementary learning systems theory and enables associative retrieval across passages that share no lexical overlap, yielding substantial gains over standard RAG baselines on multi-hop QA benchmarks. SHIMI [190] proposes a decentralized, semantic hierarchical memory index designed for scalable multi-agent reasoning, where each agent maintains a local memory shard organized by semantic similarity, and a coordination protocol enables cross-agent memory retrieval without centralizing all information in a single store. Liu et al. [191] demonstrate that fine-tuning a small LLM specifically for memory-enhanced conversation (retrieval, summarization, and persona maintenance) can transform a base model into a conversational agent with persistent memory, illustrating that the memory management pipeline itself can be a trainable component.
Experiential and reflective memory.
A substantial line of T2-aligned research focuses on memory modules that learn from experience. These tools allow a frozen agent to store, reflect on, and learn from its own output (e.g., entire trajectories), often using verbal reinforcement or self-correction, thereby building a curriculum of strategies and avoiding repeated failures without updating the core LLM’s weights [35, 192, 193]. Reflexion [35] pioneered verbal reinforcement learning: after each task attempt, the frozen agent generates a natural-language self-critique that is appended to an episodic memory buffer, and subsequent attempts condition on these reflections. The approach converts scalar reward signals into rich textual feedback without gradient updates, achieving 91% pass@1 on HumanEval and strong results on AlfWorld and HotPotQA. Retroformer [192] extends this idea by training a dedicated retrospective model that generates targeted feedback for the frozen actor, separating the “what went wrong” analysis from the “try again” generation. Think-in-Memory [193] introduces a recall-then-post-think pipeline: the agent first retrieves relevant historical thoughts from memory, then performs post-thinking to recombine and adapt them to the current context, enabling long-term coherence without expanding the context window. Agent Workflow Memory (AWM) [194] takes a complementary approach by extracting reusable workflows (structured action sequences) from an agent’s past successful trajectories and storing them in a dedicated memory module. When the agent encounters a new task, it retrieves the most relevant workflow and uses it as a template, reducing planning errors on web navigation benchmarks by 24% compared to agents without workflow memory. AWM shows that the unit of experiential memory need not be a raw trajectory or a verbal reflection; structured procedural abstractions can serve as a more efficient retrieval target. Pan et al. [195] investigate memory construction and retrieval specifically for personalized conversational agents, showing that the choice of memory granularity (individual facts vs. user-level summaries) and retrieval strategy (recency-biased vs. relevance-biased) significantly affects the agent’s ability to maintain coherent, personalized dialogue over hundreds of turns. When these accumulated experiences are distilled into reusable, composable capabilities, the result is a skill library, discussed in a dedicated paragraph below.
Structured memory (graphs, trees, and databases).
To move beyond linear text, some T2 memory tools structure information in richer forms, such as knowledge graphs, trees, or symbolic databases. The frozen agent’s outputs are used as signals to “tune” this structured tool, for example by adding new nodes, updating relationships, or writing to a database. AriGraph [196] builds an episodic knowledge graph during interaction: the agent emits observations that are parsed into entity-relation triples and merged into a persistent graph, which the agent later queries for multi-hop reasoning. ChatDB [197] externalizes memory into a relational database, translating the agent’s natural-language outputs into SQL operations (INSERT, SELECT, UPDATE) that maintain structured records across conversation turns. Rezazadeh et al. [198] propose a hierarchical tree memory that organizes conversations into nested topic clusters, enabling efficient retrieval at multiple granularity levels. Zep [199] introduces a temporal knowledge graph architecture for agent memory that explicitly models time as a first-class dimension, allowing the agent to reason about when facts were learned, how they have changed, and which are most current, a capability that flat vector stores lack. These structured representations can be more efficiently queried and reasoned over by the frozen agent, effectively externalizing complex memory management into a specialized, adaptive tool [200].
Parametric and hybrid memory architectures.
A parallel line of research explores parametric memory mechanisms that complement explicit external stores. Memory3 [201] introduces a three-tier memory hierarchy for language models: (1) model weights as implicit long-term memory, (2) an explicit memory pool of retrievable text chunks as semi-parametric memory, and (3) the context window as working memory. A lightweight “memory circuitry” learns to route information between tiers, enabling the model to offload factual knowledge to the explicit pool while reserving parametric capacity for reasoning. On language modeling benchmarks, Memory3 with a 2.4B parameter model and an external memory pool matches the perplexity of a 6.4B model without external memory, demonstrating that explicit memory can substitute for raw parameter count. Titans [202] incorporates a differentiable long-term memory module directly into the attention mechanism, maintaining a persistent memory state that is updated through gradient-based learning during inference. Unlike standard transformers that rely solely on the fixed context window, Titans can selectively store and retrieve information across arbitrarily long sequences. These parametric approaches complement the external-store paradigm (MemGPT, Generative Agents) by demonstrating that memory adaptation can also occur within the model’s computational graph, blurring the boundary between T1 (pre-trained memory modules) and T2 (agent-supervised memory adaptation).
Episodic memory as a trainable module.
Memento [23] shows that an agent’s memory system can be optimized as an external tool without any modification to the LLM planner. The system combines a frozen GPT-4.1 high-level planner with a trainable episodic case memory module. The memory stores past problem-solving trajectories, and the tool being trained is a neural Q-function that learns a case retrieval policy: which past cases to present to the frozen planner when facing a new problem.